solve
8 log2.5x=12.5

Answer by stanbon(75887)

solve
8 log2.5x=12.5
---
log(2.5x) = 12.5/8 = 1.5625
2.5x = 10^1.5625
x = 36.52
=================
Cheers,
Stan H.
=================